In how many ways can a sample (without replacement) of 4 items be selected from a population of 6 items?
Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(6,4) = (6!)/(4!*(6-4)!)

C(6,4) = (6!)/(4!*2!)

C(6,4) = (6*5*4!)/(4!*2!)

C(6,4) = (6*5)/(2!)

C(6,4) = (6*5)/(2*1)

C(6,4) = (30)/(2)

C(6,4) = 15

So there are 15 ways.